OpenStudy (lawlesscrob):

The cost of tuition at Johnson Community College is $200 per credit hour. Each student also has to pay $70 in fees. Write an equation that represents the tuition cost C for x credit hours taken. a. C = 200x + 70 c. C = 200 + 70x b. C = 70x d. C = 200x

OpenStudy (queenie):

C = 200x + 70...x is representing the number of credit hrs...do u understand ?

OpenStudy (queenie):

for example..if u had 2 credit hrs.. C = 200(2) + 7 C = 300 + 70 C = 470

OpenStudy (lawlesscrob):

yeah I think so, Thanks

OpenStudy (queenie):

typo * C = 400 + 70

OpenStudy (lawlesscrob):

Thank you